If 3 5 + 125 = 17.88, then what will be the value of 80 + 6 5 ? 13.41 B. 20.46 C. 21.66 D. 22.35











If 35 + 125 = 17.88, then what will be the value of 80 + 65 ?


















13.41 B. 20.46
C. 21.66 D. 22.35

Grade:upto college level

1 Answers

Amit Saxena
10 years ago

Answer: Option D

Explanation:

35 + 125 = 17.88

 35 + 25 x 5 = 17.88

 35 + 55 = 17.88

 85 = 17.88

 5 = 2.235

 80 + 65 = 16 x 5 + 65

   = 45 + 65

   = 105 = (10 x 2.235) = 22.35